Masses of 50g, 100g, and 200g are placed at the 30 cm, 60 cm, and 80 cm mark, respectively, of a meterstick. If the mass of the meterstick is 300g, locate the center of gravity of the system. 


1
Expert's answer
2021-07-12T12:14:43-0400
"50(50-30)+300(50-x)\\=100(60-50)+200(80-50)\\x=30\ cm"